Bob paid a total of $32,000 for his new car, including his down–payment and the interest on a loan. If the interest rate was 5%, and he made a down-payment of $5000, what was the sticker price of the car?
Answer by checkley71(8403)

32,000-5,000=27,000 which was the loan for 5% thus .05*27000=1,350 interest
32,000-1,350=30,650 which is the basic price of the car.